OpenStudy (anonymous):

Solve: 4x2 + 12x = 0

OpenStudy (saifoo.khan):

0 or -3

OpenStudy (anonymous):

\[4x^2+12x=0\] \[4x(x+3)=0\] \[x=0\] or \[x+3=0\] \[x=-3\]